### v1

# Leveraging Vision-Language Models for Visual Grounding and Analysis of Automotive UI

Abstract
Modern automotive infotainment systems require intelligent and adaptive solutions to handle frequent User Interface (UI) updates and diverse design variations. We introduce a vision-language framework for understanding and interacting with automotive infotainment systems, enabling seamless adaptation across different UI designs. To further support research in this field, we release AutomotiveUI-Bench-4K, an open-source dataset of 998 images with 4,208 annotations. Additionally, we present a synthetic data pipeline to generate training data. We fine-tune a Molmo-7B-based model using Low-Rank Adaptation (LoRa) and incorporating reasoning generated by our pipeline, along with visual grounding and evaluation capabilities. The fine-tuned Evaluative Large Action Model (ELAM) achieves strong performance on AutomotiveUI-Bench-4K (model and dataset are available on Hugging Face) and demonstrating strong cross-domain generalization, including a +5.2% improvement on ScreenSpot over the baseline model. Notably, our approach achieves 80.4% average accuracy on ScreenSpot, closely matching or even surpassing specialized models for desktop, mobile, and web, such as ShowUI, despite being trained for the infotainment domain. This research investigates how data collection and subsequent fine-tuning can lead to AI-driven progress within automotive UI understanding and interaction. The applied method is cost-efficient and fine-tuned models can be deployed on consumer-grade GPUs.
Key Words: Graphical User Interface, Large Action Model, Vision-Language Model, Automotive UI, Fine-Tuning.
ADAS
Advanced Driver-Assistance Systems
ECU
Electronic Control Unit
ELAM
Evaluative Large Action Model
GUI
Graphical User Interface
HiL
Hardware-in-the-Loop
HMI
Human-Machine Interface
LAM
Large Action Model
LLM
Large Language Model
LoRa
Low-Rank Adaptation
OEM
Original Equipment Manufacturer
SoM
Set-of-Mark
UI
User Interface
VLM
Visual Language Model
OCR
Optical Character Recognition
1 Introduction
Automotive infotainment systems are rapidly evolving, characterized by increasing complexity, dynamic interfaces, and personalization [20]. With manufacturers frequently pushing over-the-air updates and diverse designs proliferating across models, these systems demand intelligent and adaptive solutions capable of handling significant variations [36]. This constant evolution necessitates intelligent systems that can dynamically interpret the visual layout and semantic meaning of interfaces, moving beyond reliance on fixed structural assumptions. Visual Language Models offer a promising approach by integrating computer vision with natural language understanding, enabling systems to interpret visual information and user intent in a more human-like manner [17]. While these models have demonstrated success in understanding and interacting with interfaces in domains like desktop, mobile and web [3, 10], their application to the distinct environment of automotive infotainment systems remains significantly under-explored. This represents a research gap, given the unique demands and rapid evolution of in-vehicle interfaces. Applying VLMs effectively within the automotive context introduces non-trivial challenges. Infotainment Human-Machine Interfaces exhibit vast heterogeneity across car models and manufacturers, featuring custom icons, menus, and interaction paradigms [20]. Therefore, a robust model must generalize effectively across disparate screen layouts and styles while maintaining high precision in understanding on-screen elements and their context. Moreover, sophisticated interaction often requires reasoning about the HMI state, extending beyond simple object detection or element localization.
To bridge this gap, we introduce in this work a highly adaptable vision-language framework designed for understanding and interacting with automotive infotainment systems, enabling seamless adaptation across different UI designs. Our key contributions include:
  1. 1.
    The fine-tuned Molmo-based Evaluative Large Action Model ELAM-7B, optimized for automotive UI understanding, capable of processing visual input and natural language test action and evaluation instructions.
  2. 2.
    The release of AutomotiveUI-Bench-4K, an open-source dataset featuring 998 infotainment images with 4,208 annotations, hosted on Hugging Face, to serve as a benchmark and foster research in this domain.
  3. 3.
    A synthetic data generation pipeline developed to enhance model performance and generalization for small VLMs (7B or less) through parameter efficient fine-tuning.
Through empirical evaluation, we demonstrate the efficacy of our proposed framework. The fine-tuned model establishes a new performance benchmark on our AutomotiveUI-Bench-4K dataset and exhibits remarkable adaptability, significantly improving results on the cross-domain ScreenSpot task (+5.2%). Strikingly, its overall accuracy (80.4% average) competes favorably with, and in some cases surpasses, models specifically designed for narrower domains like desktop, mobile, or web (e.g., ShowUI [18]) as shown in Table 4. The robustness of the model is highlighted by its performance even with a restricted training dataset, both in terms of size and domain. It is important to note, however, that the automotive infotainment domain itself encompasses a diverse range of systems and functionalities. The following sections delve into the system architecture, the curation of our novel dataset, and the comprehensive experimental validation, collectively showcasing the significant potential of AI-driven methods for advancing automotive HMI understanding and interaction.
2 Research Scope
The validation of automotive infotainment HMI traditionally relies on established methodologies such as specification-based testing, which verifies HMI functionality against documented requirements in a black-box manner [14], and Hardware-in-the-Loop (HiL) testing, where the infotainment unit operates in a simulated vehicle environment [15, 16]. Within HiL setups, HMI validation often employs visual techniques like Optical Character Recognition (OCR) and template matching to assess the Graphical User Interface (GUI) based on captured screen images [21, 28]. However, these conventional methods, particularly their dependence on OCR and template matching for visual verification, struggle with the complexities of modern, dynamic UI. They exhibit fragility towards visual variations arising from updates, themes, or different languages [26, 5], and lack a deeper semantic understanding of the interface elements and their context [7]. The resulting high maintenance overhead and limited adaptability motivate the need for more intelligent solutions. Therefore, this research focuses on leveraging VLMs to overcome these limitations, aiming to establish a more robust, flexible, and semantically aware framework for understanding and validating automotive HMI. Notably, this work concentrates on the development and evaluation of the VLM-based framework itself and does not provide specific recommendations or methodologies for its integration into production-ready HMI/HiL testing environments.
3 Method
This chapter outlines our methodology for adapting our framework to automotive infotainment UIs. We discuss the integration of VLMs for tasks like image captioning and visual question answering, and address challenges in adapting these models for automotive UI. We introduce our approach, including synthetic data generation and fine-tuning of baseline models, to develop a specialized VLM for accurate visual grounding and validation.
3.1 VLM Integration
VLMs have emerged as a pivotal area of research, bridging the gap between visual and textual information processing. These models integrate visual capabilities with Large Language Models to handle complex tasks such as image captioning, visual question answering, and multimodal dialogue systems. Recent advancements in VLMs have led to their classification into three main categories based on their input-processing and output-generation capabilities: Vision-Language Understanding Models, Multimodal Input Text Generation Models, and Multimodal Input-Multimodal Output Models [9].
The evolution of VLMs has also seen the development of models specializing in GUI interaction. Recently the research community established the term Large Action Model (LAM) [29]. Notable comprehensive surveys [37, 22, 31] address critical research questions regarding the development of LAMs tailored for GUI tasks, and the evaluation metrics as well as benchmarks such as ScreenSpot [3], necessary to assess their effectiveness. These studies identify that a specialized VLM that is capable of accurate visual grounding is crucial to perform well in various GUI domains such as desktop, mobile, and web [37].
3.2 Domain Adaptation Challenges
As discussed in the previous section, extensive research covers LAMs and GUI agents, which focus on interactions and implicitly derive evaluative functions by inferring necessary actions. In our targeted application of validating requirements via test cases defined in natural language, it is necessary to explicitly evaluate expected outcomes. Small fine-tuned models (7B or less) such as TinyClick [24] or UGround [10] gathered UI data from mobile, desktop, and web applications to fine-tune baseline models such as Qwen2-VL [30] for visual grounding. This raises the following challenges:
  1. 1.
    Adapt VLMs for the domain of automotive UI: Modern infotainment systems contain many different UI designs depending on brand, platform, mode of driving and display hardware properties. Icons pose a particular challenge as there is little overlap to generic UIs from desktop, mobile, or web applications.
  2. 2.
    Assert evaluation capability: Prior work does not consider explicit evaluation. In order to evaluate requirements or specifications it is necessary to prompt expected results which are visually grounded and evaluated.
  3. 3.
    Ensure onsite deployment and control of models:
    • Original Equipment Manufacturers often have strict information security requirements which makes leveraging proprietary models such as OpenAI GPT-4o or Google Gemini Pro 2.0 challenging due to concerns about data privacy and potential leakage of sensitive information to external servers.
    • Another challenge arises from fine-tuning such models. OpenAI and Google provide fine-tuning services for training and deployment. Details of the training process and the fine-tuned model itself remain closed source.
    • The discontinuation of closed source models entail challenging validation and prompt adaptation iterations in order to ensure the intended functionality.
    • Running model on local servers or even consumer-grade GPUs is highly desirable and necessary to assure a cost-effective solution.
To address these challenges, we introduce ELAM-7B, that goes beyond existing LAMs by integrating evaluation capabilities. This allows the model to conduct test cases. The model is trained with synthetic data generated by our accompanying pipeline.
3.3 Synthetic Data Pipeline
Object-level captioning for infotainment screens presents significant challenges due to the necessity of incorporating functional, positional, and visual attributes to uniquely identify each element. As previously mentioned, differentiating between Test Actions and Expected Results is required. Both classes necessitate object-level captions that describe either the command initiating element actuation or a statement that shall be evaluated for the current screen. Furthermore, for Expected Results, a binary status (passed or failed) must be assigned, reflecting the validity of the expectation within the image. Empirically, human annotators exhibit a tendency to under-represent failed expectations, likely due to the cognitive burden of formulating incorrect expectations for a displayed infotainment screenshot. Therefore, a hybrid approach, leveraging human expertise for UI element selection via bounding boxes combined with automated caption generation by large teacher models, (e.g., GPT-4o, Gemini 2.0 Flash Thinking), offers a practical balance between annotation quality and time investment. Automated UI element extraction methods like OmniParser V1.5 [19] and a custom YOLOv8 [13] fine-tuned to automotive UI images were found to be insufficiently reliable, frequently resulting in misaligned bounding boxes that negatively impacted the correctness of generated captions.
The following synthetic data pipeline (Fig. 1) was used to train ELAM-7B. Bounding boxes from the dataset were prompted using the Set-of-Mark (SoM) technique [35]. Inspired from UGround [10], we used a bounding box pointed to by an arrow. For Test Actions, prompts were designed to short yet comprehensive instructions that unambiguously map to a single UI element (Fig. 2(a)). Prior to generating each test action, the model was instructed to provide a reasoning detailing the process of identifying the target UI element, including its semantic, positional, and visual characteristics, as well as any relevant relationships to other elements, such as the current context or list headings (Fig. 2).
Figure 1: Synthetic data generation pipeline
(a) Synthetic Test Action
(b) Synthetic Expected Result
[…]
To improve the generation of Expected Results, the process was modified to include a prior successful test action leading to the current menu or screen state. Specifically, the model first generated a prior test action, followed by reasoning for the expectation and its “passed/failed” conclusion. Initially, this approach exhibited the same bias as human annotators and generated mostly passed validations. This issue was mitigated by explicitly prompting for evaluations that entail a failed result (Fig. 2(b)). After manually reviewing a small set of samples for GPT-4o, Gemini 2.0 Flash Thinking and Claude 3.5 Sonnet, we decided to use Gemini (gemini-2.0-flash-thinking-exp-01-21) to generate the data for the training procedure. The detailed prompts used for generating the training datasets are displayed in Sections A.1.1, A.1.2 and A.1.3.
To mitigate potential overfitting on the specific linguistic style of the teacher model, we rephrased the text for reasoning, test actions, and expectations using a smaller model (gpt-4o-mini-2024-07-18). This rephrasing step is crucial for ensuring robustness and generalizability when processing inputs from real-world test case data. The teacher model, configured with a temperature of zero to minimize hallucinations, exhibited a tendency towards repetitive use of verbs and sentence structures, necessitating this diversification strategy. The contribution of rephrasing is explored in Table 5.
3.4 Model Fine-Tuning
For efficient training of baseline models, we utilized LoRa [12]. LoRa offers benefits such as accelerated training, reduced training data needs, smaller storage footprint for model artifacts (adapters), lower VRAM requirements, and the potential for dynamic adapter swapping in deployment, eliminating the need to load the complete model.
To establish a suitable baseline model for our fine-tuning and experiments, we evaluated four existing VLMs: TinyClick (0.27B) [24], Molmo-7B-D-0924 [4], UGround-V1-7B (Qwen2-VL-based) [10], and InternVL2_5-8B [2]. TinyClick and UGround are fine-tuned versions of Florence-2-Base [34] and Qwen2-VL [30], respectively. These models are specifically designed to process UI images, interpret actions, and generate corresponding point coordinates. However, their broader capabilities beyond this visual grounding task are limited. They frequently output points even when the prompt requests different information. TinyClick’s relatively small size further restricts its potential as a robust foundation for our work. In contrast, Molmo and InternVL2.5 are general-purpose VLMs. Notably, they were also trained on datasets including GUI data from mobile and web applications. While InternVL2.5 can generate bounding boxes, Molmo is limited to point generation. When evaluated on our proposed AutomotiveUI-Bench-4k dataset, Molmo-7B-D-0924 achieved the best performance (Table 3) among the considered models and was therefore chosen as the foundation of ELAM.
We generated a dataset of 17,708 annotations for 6,230 images using the synthetic data pipeline described in Section 3.3. It comprises 5,952 Test Actions, 6,190 passed, and 5,566 failed evaluations with corresponding Expected Results. The images in our training datset were captured using 4k cameras, either from our vehicle fleet or with the generous consent of local car dealerships. This dataset was subsequently utilized for LoRa fine-tuning of the Molmo-7B-D-0924 checkpoint. The prompt templates used for training, evaluation, and inference are described in Table 1.
Training was conducted for 2 epochs with a global batch size of 128, achieved using a local batch size of 8 across two NVIDIA H100 80 GB PCIe GPUs. We employed a learning rate of 1e-4 and incorporated LoRa to enhance parameter efficiency, applying it to all linear layers within Molmo.
Table 1: Prompt and response templates for training and inference of ELAM as Python f-strings
Prompt Template - Test Action
{python} prompt_template_test_action = f””” Identify and point to the UI element that corresponds to this test action: test_action. ”””
Response Template - Test Action
{python} response_template_test_action = f””” reasoning ¡point x=”center_x:.1f” y=”center_y:.1f” alt=”test_action”¿ test_action¡/point¿ ”””
Prompt Template - Expected Result
{python} prompt_template_expected_result = f””” Evaluate this statement about the image: ’expectation’ Think step by step, conclude whether the evaluation is ’PASSED’ or ’FAILED’ and point to the UI element that corresponds to this evaluation. ”””
Response Template - Expected Result
{python} response_template_expected_result = f””” reasoning Conclusion: evaluation_result ¡point x=”center_x:.1f” y=”center_y:.1f” alt=”expectation”¿ expectation evaluation_result¡/point¿ ”””
4 Open Source Infotainment Validation Dataset: AutomotiveUI-Bench-4K
Given that multiple datasets [25, 33, 32, 6, 38, 1, 8] already exist in the fields of desktop, mobile, and web, we identify a domain gap in scientific research for benchmarking LAMs for automotive UI. To fill this gap, we introduce and release an infotainment UI dataset named AutomotiveUI-Bench-4K. This dataset comprises 998 images with 4,208 annotations from 15 brands such as Audi, BMW, Ford, Porsche, and Tesla among others, thus ensuring a diverse variety of modern vehicle UI, meticulously crafted by experienced professional software testers specializing in HMI evaluation. In addition to images from automotive OEMs, the dataset also incorporates screenshots of Apple CarPlay and Google Android Auto. Images from OEMs were captured in 4K resolution using cameras and sourced from either our vehicle fleet or through collaboration with local car dealerships. To ensure consistent image perspective and focus on the relevant user interface elements, the OEM images were rectified and cropped to the active display area. Android Auto and CarPlay screenshots were directly exported. The infotainment systems featured in this dataset are exclusively touchscreen-based. While some systems include auxiliary physical buttons, these were not considered in our analysis due to their primarily static functionality. Although the dataset primarily comprises systems from recent model years, the system representing the earliest model year is a 2018 Volkswagen Tiguan. Languages that were included are German and English. Instructions and evaluation requests are written in English. German text was either translated or directly quoted. Real-world test engineering documentation often exhibits a pattern of short sentences and occasional grammatical deviations. Therefore, to create a dataset that realistically represents this domain, it is crucial to preserve these stylistic features in the preprocessing steps. Table 2 shows further information about the data distribution. The actions and results are unrelated, since test actions can lead to different screens and we do not consider trajectories. There are two annotation classes in the dataset:
  1. 1.
    Test Action: Describes an interaction with a single UI element as an imperative sentence (”set A/C to max” in Fig. 3).
  2. 2.
    Expected Result: Represents a testable expectation, defined per image, focusing on the required state or behavior of a particular UI element within that image. This expectation is evaluated to determine a passed/failed outcome. (”Passenger’s climate zone is synced to driver” in Fig. 3)
Table 2: Label distribution in AutomotiveUI-Bench-4K
Category Subcategory Total EN DE
Images - 998 454 544
Annotations - 4,208 1,988 2,220
Test Action - 2,269 1,059 1,210
Expected Result Total 1,939 929 1,010
Expected Result Passed 1,375 662 713
Expected Result Failed 564 267 297
Figure 3: Example for a test action (red) “set A/C to max” and for an expected result (green) “Passenger’s climate zone is synced to driver” (Passed)
5 Experiments
5.1 Evaluating Performance with AutomotiveUI-Bench-4K
The performance of various baseline models and fine-tuned models are shown in Table 3. The numeric values represent accuracy in percent. $TA$ and $ER$ represent data samples for classes Test Action and Expected Results, respectively. Visual grounding scores are indicated by the index $vg$. Evaluations pertaining to predicted outcomes are denoted by the index $evl$. Accuracy was assessed by analyzing the generated point or centroid of the bounding box and verifying its containment within the annotated ground truth box. To assess language-specific performance, the image dataset was stratified into two distinct categories based on the language (English or German) represented in the images. When a generated evaluation could not be parsed from the model’s response, it was assigned the inverse of the ground truth value. All evaluation experiments for AutomotiveUI-Bench-4K were conducted using a NVIDIA RTX 4090 GPU with 24 GB VRAM.
The findings suggest that models within the 7B parameter range exhibit no significant performance degradation when processing images with predominantly German text. A significant challenge for most baseline models was the generation of consistently parseable output. For example, InternVL2.5 had good evaluation performance but rarely generated bounding boxes. TinyClick and UGround were only evaluated for their grounding performance as they do not have evaluative capabilities. To explore the possibility of faster inference through task distribution across models, we experimented with LoRa-fine-tuning TinyClick (LAM-270M), due to its size of only 0.27 billion parameters. This approach, however, did not produce performance on par with Molmo-7B as shown in Table 3, and we therefore abandoned further investigation in this direction.
Table 3: Evaluation on AutomotiveUI-Bench-4K
Model $\downarrow\!TA_{vg}$ $TA_{vg}^{DE}$ $TA_{vg}^{EN}$ $ER_{vg}$ $ER_{vg}^{DE}$ $ER_{vg}^{EN}$ $ER_{evl}$ $ER_{evl}^{DE}$ $ER_{evl}^{EN}$
InternVL2.5-8B* [2] 26.6 26.1 27.0 5.7 6.3 5.1 64.8 60.4 69.0
TinyClick [24] 61.0 54.6 67.8 53.3 47.3 59.2 - - -
UGround-V1-7B (Qwen2-VL) [10] 69.4 68.9 69.9 55.0 54.4 55.7 - - -
Molmo-7B-D-0924 [4] 71.3 70.9 71.8 71.4 69.8 72.9 66.9 67.8 66.0
LAM-270M (TinyClick) 73.9 66.9 81.1 59.9 54.6 65.1 - - -
ELAM-7B (Molmo) 87.6 87.5 87.6 77.5 77.0 77.9 78.2 78.5 77.8
    Since the InternVL2.5 series model natively supports bounding boxes, these were used for evaluation instead of points.
5.2 Evaluating Generalizability with ScreenSpot
To determine if our fine-tuning resulted in overfitting to the automotive UI domain, we applied ELAM-7B to the ScreenSpot dataset [3] comparing it to SeeClick [3], ShowUI [18], OS-Atlas-Base-7B [33], Molmo-7B-D-0924/Molmo-72B-0924 [4], and Qwen2-VL-based UGround-V1-7B [10] in Table 4. Our fine-tuning procedure did not break the generalizability of Molmo. The average score per category was improved by 5.2% compared to the baseline (Molmo-7B-D-0924) and by 2.2% compared to Molmo-72B.
Table 4: Evaluation results on ScreenSpot for selected models taken from [11]
[…]
SeeClick [3] 53.4 78.0 52.0 72.2 30.0 55.7 32.5
ShowUI-2B [18] 75.1 92.3 75.5 76.3 61.1 81.7 63.6
Molmo-7B-D-0924 [4] 75.2 85.4 69.0 79.4 70.7 81.3 65.5
UGround-V1-2B (Qwen2-VL) [10] 77.7 89.4 72.0 88.7 65.7 81.3 68.9
Molmo-72B-0924 [4] 78.6 92.7 79.5 86.1 64.3 83.0 66.0
ELAM-7B (Molmo) 80.8 94.5 79.5 89.2 70.7 85.7 65.0
OS-Atlas-Base-7B [33] 81.0 93.0 72.9 91.8 62.9 90.9 74.3
UGround-V1-7B (Qwen2-VL) [10] 86.3 93.0 79.9 93.8 76.4 90.9 84.0
5.3 Ablation Study
We performed ablation studies to evaluate the contributions of simple reasoning and rephrasing modules integrated into our synthetic data generation pipeline. Training runs were conducted using distinct preprocessing strategies, the configurations of which are summarized in Table 5. The prompts from Table 1 were used for experiments that incorporate reasoning, whereas the phrase “Think step by step, conclude” was replaced by “Determine” for runs that did not utilize reasoning. A LoRa-rank of 64 was chosen to fine-tune all linear layers for the ablation experiments. Our findings demonstrate that the inclusion of reasoning significantly enhances performance, as evidenced by improvements in both visual grounding and evaluation accuracy. This observation highlights the critical role of reasoning for robust generalization to real-world data in potential future deployments.
To optimize the LoRa configuration, we conducted a separate ablation study focusing on the rank parameter $r$, the results of which are presented in Table 6. The data was preprocessed utilizing reasoning and rephrasing. This analysis determined that a rank of 64 yielded optimal performance. In all LoRa-based experiments, the alpha parameter was set to $r$, and the LoRa-dropout rate was fixed at $0.05$.
Table 5: Prompting ablations
Settings $TA_{vg}$ $ER_{vg}$ $ER_{evl}$
Baseline 81.7 73.5 73.5
Rephrasing 83.9 73.7 71.1
Reasoning 86.4 76.2 77.0
Reas.+Rephr. 87.4 77.2 78.6
Table 6: LoRa ablations
Rank $TA_{vg}$ $ER_{vg}$ $ER_{evl}$
16 84.6 76.1 77.7
32 85.0 76.4 77.1
64 87.4 77.2 78.6
128 87.1 77.2 77.6
5.4 Analysis of Utterance Embedding Space vs. VLM Performance
To investigate the relationship between semantic representation of user utterances and VLM performance, particularly within car application subdomains, we analyzed two distinct task types: Test Actions and Expected Results as outlined in Section 4. For both types, we generated text embeddings of the utterances using the nomic-embed-text-v1.5***nomic-embed-text-v1.5 is an improved variant of the nomic-embed-text-v1[23] model. It utilizes Matryoshka Representation Learning for flexible dimensionality reduction with minimal loss while supporting up to 8,192 tokens. The model is optimized for search, clustering, and classification in production. model and visualized their structure using t-SNE dimensionality reduction [27]. Unsupervised k-Means clustering ($k=8$) was applied to the high-dimensional embeddings to identify potential thematic groups, with cluster labels generated post-hoc via LLM (GPT-4o) interpretation of sampled utterances. Crucially, we enhanced the t-SNE projections with a heatmap that visualizes the local failure rate for each designated grounding for Test Actions and both verification and grounding for Expected Results. This failure rate was calculated per grid cell based on evaluation results, with red indicating high failure density and white indicating low failure density. This approach allows us to visualize how VLM performance, for both the base Molmo-7B and ELAM-7B models, varies across the fixed semantic landscape defined by the text embeddings.
5.4.1 Test Action Utterances: Grounding Performance
Baseline Molmo-7B-D-0924:
The t-SNE plot (Fig. 4) of Test Action embeddings for the base model shows clusters with LLM-suggested labels indicating semantic overlap (e.g., ’Navigation Control’ across multiple clusters). The grounding failure heatmap reveals that errors are distributed across the embedding space, not perfectly aligning with cluster boundaries. Notable high-failure regions overlap with clusters tentatively labeled Climate Control, Audio and Display Settings and User Interface Control, Device Settings. A distinct cluster associated with Navigation Control, User Interface Control resides in a predominantly low-failure region, suggesting these utterance types are generally grounded successfully.
Figure 4: t-SNE plot of the base Molmo-7B model for the Test Action task
ELAM-7B:
Using the same text embeddings, the fine-tuned model’s performance is visualized through its distinct failure heatmap (Fig. 5). When comparing Fig. 5 to Fig. 4, a noticeable reduction in both the extent and intensity of high-failure regions across the embedding space can be observed, supporting the results summarized in Table 3. Several clusters show significant improvement. For instance, the most distinct Cluster 3: Navigation Control, User Interface Control achieved a failure rate of close to zero. The dispersed Cluster 1: User Interface Control, Communication Management, Cluster 4: Climate Control, Vehicle Settings and Cluster 5: Climate Control, Vehicle Settings now predominantly reside in low-failure regions after fine-tuning, indicating that the model has learned to handle these automotive domain specific utterance types more reliably. However, some high-failure samples still persist at the boundaries of these clusters, suggesting that certain semantic regions or utterance formulations remain challenging even after fine-tuning.
Figure 5: t-SNE plot of the ELAM-7B model for the Test Action task
5.4.2 Expected Result Utterances: Verification and Grounding Performance
Baseline Molmo-7B-D-0924:
The Expected Result utterances, representing a more complex task involving visual grounding and subsequent evaluation, also form clusters in the embedding space with themes like ’Settings’, ’Navigation’, and ’Display’. The base Molmo-7B failure heatmap (Fig. 6) for this task shows widespread and often intense high-failure regions, particularly concentrated in areas associated with User Interface Settings, Search Functionality and Audio Control, Radio Management.
Figure 6: t-SNE visualizations of the base Molmo-7B model for two Expected Result evaluation tasks: (a) Visual Grounding and (b) Evaluation.
ELAM-7B:
The failure heatmap (Fig. 7) for the fine-tuned model on the Expected Result tasks shows that high-failure regions persist across many parts of the embedding space. While a visual comparison with the base Molmo model might suggest a subtle reduction in the overall failure density, the improvement appears less pronounced than observed for the Test Action task.
Figure 7: t-SNE visualizations of the ELAM-7B model for two Expected Result evaluation tasks: (a) Visual Grounding and (b) Evaluation.
Fine-tuning provided only modest improvements for the more complex Expected Result verification and grounding task. The widespread persistence of high-failure regions suggests that accurately interpreting and verifying diverse state descriptions based on visual evidence remains a significant challenge.
5.4.3 Summary
The t-SNE visualization of a fixed text embedding space overlaid with task-specific VLM failure rates provided valuable insights by decoupling utterance semantics from model performance. The analysis indicated tangible benefits of fine-tuning for more direct tasks such as the Test Action grounding, while suggesting that more complex tasks like the Expected Result verification still present significant challenges. Additionally, the t-SNE plots highlighted variations in performance across different sub-domains. The model performed well in grounding UI elements commonly found in desktop, mobile, and web environments but showed reduced effectiveness in automotive-specific functionalities such as Advanced Driver-Assistance Systems (ADAS) or climate control, for instance, distinguishing between adjusting the driver’s and passenger’s temperatures. This underscores the general need for further domain-specific data and training strategies.
6 Conclusion
Our research demonstrates a clear improvement in visual grounding performance using our LoRa fine-tuned model with a synthetic generation pipeline for its training process. The model outperformed its baseline on both AutomotiveUI-Bench-4K and ScreenSpot, highlighting our approach as an effective strategy for enhancing visual grounding abilities with limited resources. While our findings are encouraging, it is important to acknowledge certain limitations of our current study.
Firstly, while the synthetic data pipeline proved effective, the inherent gap between synthetic and real-world data introduces a potential limitation of our findings to entirely unconstrained scenarios. Further investigation is needed to assess performance on a wider range of real-world datasets beyond AutomotiveUI-Bench-4K, and across more diverse visual grounding tasks. Additionally, the scope of our evaluation, while demonstrating improvement, primarily focused on quantitative metrics. A deeper qualitative analysis of the model’s performance, especially on automotive-specific parts of infotainment systems, would provide richer insights into the nuanced strengths and weaknesses of our approach.
Looking ahead, there are several promising avenues for potential improvements. Future work could focus on developing more advanced and general evaluation methodologies that integrate linguistic and grammatical analysis to better understand how subtle syntactic and semantic nuances affect model performance. This may include incorporating techniques like dependency parsing and syntactic analysis to provide a more comprehensive picture of the interplay between language and visual grounding.
Moreover, there is significant potential for enhancing the capabilities of the visual encoders within VLM architectures. Although current pre-trained encoders are effective, they may overlook certain domain-specific features and high-resolution details. Future improvements might explore the use of multi-encoder strategies, adaptive encoder selection, and more data-efficient fine-tuning techniques to overcome these limitations and achieve more robust and generalized performance across diverse real-world scenarios.
Overall, our findings point toward a holistic strategy for advancing VLM technologies, one that not only refines visual grounding through innovative training pipelines, but also embraces broader evaluation frameworks and encoder enhancements. These integrated improvements hold the promise of pushing the boundaries of current models toward greater real-world applicability.
Acknowledgement
This work was partially supported by German BMBF within the research project MANNHEIM-KI4BoardNet. (https://www.elektronikforschung.de/projekte/mannheim-ki4boardnet).
Appendix A Appendix
A.1 Prompts in Synthetic Data Pipeline
A.1.1 Test Action Template
Write test actions for the UI of automotive infotainment software.
These actions must be declarative, concise, and tailored to verify the softwares functionality accurately.
Test actions should cover marked user interface elements and interactions relevant to the infotainment system.
# Steps
1. Understand UI Element:
- Identify the position of the UI element marked by a <color> <marker_type>.
- Identify the semantic meaning based on its text or icon.
- Identify any parent elements crucial for semantic or functional meaning.
- Identify any related elements crucial for semantic or functional meaning (e.g., text corresponding to a checkbox or switch).
2. Reasoning if UI Element is interactive:
- Determine why the element is interactive or not. Grayed out elements could either mean that they are just disabled or not interactive at all.
- If it is not interactive set the utterance to none”.
3. Specify Test Action Utterance:
- Write a deterministic and declarative action simulating a user interaction with the UI element.
- Use unique identifiers for the UI element, and mention a parent element if necessary.
- Do not use the <color> <marker_type> in the test action utterance.
- Use verbs like tap, click, enter, open, enable, disable, activate, deactivate, select, press, collapse, navigate, cancel, refresh,
- Try to choose the verb that is most applicable for the type of UI Element (e.g., enable for switch, choose for radio buttons, open for submenu etc.)
# Required Output Structure
REASONING:
1. [First step in thinking process]
2. [Second step in thinking process]
[Continue with numbered steps as needed]
UTTERANCE:
[Describing the test action utterance as a single sentence without using the <color> <marker_type>]
A.1.2 Expected Result Template “Passed”
As a test engineer for automotive infotainment systems, formulate result evaluations based on the current screen and determine if the test has passed or failed.
# Steps
0. Understand the UI Element:
- Analyze the current context and active menu of the infotainment system.
- Identify the position of the UI element marked by a <color> <marker_type>.
- Determine the semantic meaning based on its text or icon.
- Identify any parent elements crucial for semantic or functional meaning.
- Identify any related elements crucial for semantic or functional meaning (e.g., text corresponding to a checkbox or switch).
1. Performed Test Action:
- Think of a possible test action which was done and can be evaluation with the element marked by the <color> <marker_type>.
2. Reasoning:
- Conduct reasoning for reaching the result evaluation by examining UI semantics with step-by-step thinking.
3. Determine Evaluation/Expected Result:
- Provide a short and general description of the expected result of the test action that led to the current screen or highlighted UI element.
- The expected result must be based solely on the current screen, not on previous or next screens.
- Include presence, color, positional, semantic, state, and visual information as needed. Do not include the <color> <marker_type>.
- The expected result can also just check the presence of the marked UI element
4. Incorporate Evaluation:
- Assess why the expected result is met or not. Conclude with FAILED or PASSED.”
# Critical Rules
1. Never reference <color> <marker_type> in test action or expected result
2. Evaluate only current screen state
3. No assumptions about previous/future states
4. Use objective, verifiable statements
5. Document all reasoning steps
6. Provide clear pass/fail criteria
7. Valid evaluations include checking the presence, visibility, position, and properties of the UI element.
# Required Output Structure
TEST ACTION:
[Single sentence describing the specific test action performed]
REASONING:
1. [First step in thinking process]
2. [Second step in thinking process]
[Continue with numbered steps as needed]
EXPECTED RESULT:
[Describing the evaluated result as a single sentence without using the <color> <marker_type>]
CONCLUSION:
[PASSED/FAILED]
A.1.3 Expected Result Template “Failed”
As a test engineer for automotive infotainment systems, formulate result evaluations based on the current screen and determine if the test has passed or failed.
# Steps
1. Understand the UI Element:
- Analyze the current context and active menu of the infotainment system.
- Identify the position of the UI element marked by a <color> <marker_type>.
- Determine the semantic meaning based on its text or icon.
- Identify any parent elements crucial for semantic or functional meaning.
- Identify any related elements crucial for semantic or functional meaning (e.g., text corresponding to a checkbox or switch).
2. Determine Evaluation/Expected Result that is wrong for the current screen:
- Provide a short and general description of the failed expected result of the test action that led to the current screen or highlighted UI element.
- You should think of an expectation that is wrong or not in the screen.
- The expected result must be based solely on the current screen, not on previous or next screens.
- Include absence, different color, wrong positional, semantic, and wrong state information as needed. Do not include the <color> <marker_type>.
- The expected result can also just check the absence of the marked UI element or if the screen shows the wrong context menue.
3. Reasoning:
- Conduct reasoning for reaching the result evaluation by examining UI semantics with step-by-step thinking.
4. Incorporate Evaluation:
- Assess why the expected result is met or not. Conclude with FAILED or PASSED.”
# Critical Rules
1. Never reference <color> <marker_type> in test action or expected result
2. The Expectation must be generated for elements within the <color> <marker_type>
3. No assumptions about previous/future states
4. Use objective, verifiable statements
5. Document all reasoning steps
6. Provide clear pass/fail criteria
7. Valid evaluations include checking the presence, visibility, position, and properties of the UI element
# Required Output Structure
REASONING:
1. [First step in thinking process]
2. [Second step in thinking process]
[Continue with numbered steps as needed]
EXPECTED RESULT:
[Describing the evaluated result as a single sentence without using the <color> <marker_type>]
CONCLUSION:
[PASSED/FAILED]